### v1

# Principal-Agent Multitasking: the Uniformity of Optimal Contracts and its Efficient Learning via Instrumental Regression

Abstract
This work studies the multitasking principal-agent problem. I first show a “uniformity” result. Specifically, when the tasks are perfect substitutes, and the agent’s cost function is homogeneous to a certain degree, then the optimal contract only depends on the marginal utility of each task and the degree of homogeneity. I then study a setting where the marginal utility of each task is unknown so that the optimal contract must be learned or estimated with observational data. I identify this problem as a regression problem with measurement errors and observe that this problem can be cast as an instrumental regression problem. The current works observe that both the contract and the repeated observations (when available) can act as valid instrumental variables, and propose using the generalized method of moments estimator to compute an approximately optimal contract from offline data. I also study an online setting and show how the optimal contract can be efficiently learned in an online fashion using the two estimators. Here the principal faces an exploration-exploitation tradeoff: she must experiment with new contracts and observe their outcome whilst at the same time ensuring her experimentations are not deviating too much from the optimal contract. This work shows when repeated observations are available and agents are sufficiently “diverse", the principal can achieve a very low $\widetilde{O}(d)$ cumulative utility loss, even with a “pure exploitation" algorithm.
1 Introduction
A firm’s owner (the principal) hires a manager (the agent) to manage the firm. The manager is responsible for several tasks, for example, managing personnel, maintaining a healthy work environment, planning and executing activities, etc. The manager can allocate a certain effort to each task, however, this effort is unobservable and unverifiable to the owner of the firm. Instead, the owner can observe a noisy signal, indicating the performance of each task. The principal can also observe the total revenue generated by the firm under the manager’s management, which will depend on the effort the manager puts into the tasks. The owner of the firm must incentivize the manager to exert effort through contracts. Since the exact effort is unobservable, the contract can only depend on the noisy signals associated with each task. Such is the basic model for the multitask principal-agent problem, which was first studied in the seminal work [1].
It often happens that the principal contracts with more than one agent. For example, in franchising scenarios, the principal is the franchiser and the agents are franchisees ([2]). The franchiser must design contracts to incentivize each franchisee to exert the appropriate amount of effort to generate revenue. Another example is e-commerce platforms, such as YouTube. YouTube is a video-sharing platform where content creators earn revenue through advertisements. According to this documentation ([3]), YouTube content creators earn 55% of the total advertisement revenue, while the remaining 45% goes to YouTube. While intuition might suggest that the principal should tailor her contracts when facing different agents to maximize her utility, in the above two scenarios, it is often observed that the contract is uniform across different agents. This leads to our first research question.
How can the uniformity of the contract terms be justified in (multitask) principal-agent problems?
As an answer to this question, I show that when the agent’s cost function is homogeneous to a certain degree and the principal’s utility takes a linear form across tasks (i.e., tasks are perfect substitutes), then the optimal contract is indeed uniform for different agents. In particular, the optimal contract only depends on the marginal utility of each task and the degree of homogeneity of the agent’s cost function. The implications of this result are practically very attractive. It implies that there is no need for the principal to discriminate against agents in an attempt to maximize her own utility, i.e., the optimal contract is fair. The result here follows closely in spirit to a result in the important work by [2]; the difference is that in their work they study a single task setting with double moral hazard while this work studies a multitask setting.
The above result shows that in the optimal contract, the principal should transfer more utility to the agent for more “important” tasks, which aligns with intuition. However, it may be the case that the exact marginal utility of each task is unknown. In this case, the principal must learn this value with observational data in order to design the optimal contract. This leads to our second research question.
How can the principal learn the optimal contract when the “importance” of each task is unknown?
Assuming again that the principal’s utility takes a linear form across tasks, estimating the marginal utilities of each task is essentially a linear regression problem, where the total revenue is the response variable, and the observed signals of each task are the covariates. However, since the observed signals are noisy measurements of the true effort, the regression problem has a measurement error problem ([4]). This work identifies the contract as a valid instrumental variable and proposes estimation and learning algorithms based on the generalized method of moments (GMM) estimator ([5]). Moreover, when repeated observations are available, the GMM estimator based on repeated observations achieves even faster convergence when the agents are “diverse”.
It seems important to point out that a problem setting very similar to the current work was studied in the (fairly) recent work [6]. Their work studied optimization algorithms in Stackelberg games from revealed preferences. As a special case, they gave a discussion on how their algorithm can be applied to the multitask principal-agent problem (though their work did not explicitly identify their problem as such). The problem of learning the optimal contract is also studied in this work, however, the approach is fundamentally different from theirs. This is because the current work also identified a “uniformity” result, which seemed to escape their analysis. The problem of estimating and learning the optimal contract in this work uses instrumental regression with the generalized method of moments method and is based on this “uniformity” result, whereas in their work the problem is solved via a two-stage optimization algorithm.
2 Related Work
Contract Theory
Contract theory studies a principal-agent problem where the principal must incentivize the agent to exert effort through contracts. Some early important works in the economics community include [7, 8, 1]. The work by [1] studies the multitask principal-agent problem, which serves as the starting point of the current work. Recently, computational aspects of contract design have been studied by many works in the computer science community. For example, [9, 10, 11] study combinatorial algorithms in contract design; [12, 13, 14] study learning algorithms for contract design.
Instrumental Regression and Measurement Error Models
Measurement error model is a well-studied problem in econometrics and statistics, for a textbook treatment the reader is referred to [4]. The instrumental regression and generalized method of moments (GMM) estimators (again for a textbook treatment see [5]) are popular methods to address endogeneity issues in models, including the measurement error model. Some recent works in computer science also notice endogeneity issues in learning/classification problems and apply the instrumental regression method in the context of strategic classification ([15]), online learning/bandits ([16]), reinforcement learning ([17]), etc.
Bandits and Exploration-Exploitation Tradeoff
Bandit problems are a type of online learning problem with partial feedback. Typically, bandit algorithms need to balance exploration (experimenting with unknown actions) and exploitation (choosing actions whose reward is estimated to be higher). Research on bandit problems is too broad to cover here, but for an overview see [18]. Recently, a number of papers have sought to understand the performance of the pure exploitation (i.e., greedy) algorithm in bandit problems. In particular, for linear bandits, the success of the greedy algorithm is explained under the framework of smoothed analysis [19, 20, 21].
3 Multitask Principal-Agent: Problem Description and the Uniformity Result
I study a multitasking principal-agent problem. The agent can exert effort in $d$ dimensions, where each dimension can be interpreted as a task. Denote effort of the agent by the vector $w\in(\mathbb{R}^{d})^{+}$, and assume that the agent incurs a private cost $c(w)\in\mathbb{R}^{+}$ when exerting effort $w$. The exact effort is unobservable and unverifiable to the principal, but a signal $x\in\mathbb{R}^{d}$ is generated and observed by the principal. The signal is assumed to satisfy
$\mathbb{E}[x|w]=w.$
The signal can be interpreted as a noisy measurement of the agent’s true effort.
The agent’s effort gives a private benefit $y$ to the principal, and it is assumed this benefit depends on the agents true efforts through the linear model:
$\mathbb{E}[y|w]=\langle\theta,w\rangle.$
Here $\theta\in(\mathbb{R}^{d})^{+}$. The value $\theta_{i}$ can be interpreted as the marginal utility of task $i$.
The principal incentivizes the agent to exert effort through a contract $\beta\in\mathbb{R}^{d}$. Under this contract, the monetary transfer to the agent is $\langle\beta,x\rangle$ when the realized signal is $x$. The agent then takes the action $w=w(\beta)$ which maximizes his expected utility under the contract $\beta$:
$w=\arg\max_{w^{\prime}}\langle\beta,w^{\prime}\rangle-c(w^{\prime}).$
Then, under the contract $\beta$, assuming the agent takes his best action $w$, the principal’s expected utility is
$\langle\theta,w\rangle-\langle\beta,w\rangle,$
i.e., the expected private benefit minus the expected payment to the agent. The optimal contract is the contract maximizing the principal’s expected utility.
Definition 1.
A function $c$ is homogeneous of degree $k$, if for any $\rho>0$ and $w\in(\mathbb{R}^{d})^{+}$, $c(\rho w)=\rho^{k}c(w)$.
The parameter $k$ is also known as the return to scale parameter.
Definition 2.
The cost function $c$ is said to satisfy the Inada conditions if the following holds.
  1. 1.
    $c$ is convex on its domain.
  2. 2.
    $\lim_{w_{i}\rightarrow 0}\partial c(w)/\partial w_{i}=0$.
  3. 3.
    $\lim_{w_{i}\rightarrow+\infty}\partial c(w)/\partial w_{i}=+\infty$.
The Inada condition is fairly standard. It essentially means the marginal cost goes from 00 to +$\infty$+ ∞ as $w_{i}$ goes from 00 to +$\infty$+ ∞. Throughout the following, the following assumption will be imposed on the cost function of the agent(s).
Assumption 1.
The cost function of the agent is homogeneous of degree $k$, and satisfies the Inada conditions.
3.1 The Uniformity Result
This subsection shows the optimal contract depends on the shape of $c(\cdot)$ only through the degree of homogeneity.
Theorem 1.
The principal’s optimal contract is $\beta=\theta/k$.
Proof.
When the posted contract is $\beta$, the agent’s expected utility when taking action $w$ is
$\langle\beta,w\rangle-c(w).$
The first-order condition gives:
$\beta=\nabla c(w).$
The first-order condition must bind under 1.
The principal’s expected utility, written as a function of $w$, is then
$\langle\theta,w\rangle-\langle\beta,w\rangle=\langle\theta,w\rangle-\langle\nabla c(w),w\rangle=\langle\theta,w\rangle-kc(w).$
Note the above is a concave function in $w$. The first-order condition gives
$\theta=k\nabla c(w).$
But also recall at the agent’s optimum $\beta=\nabla c(w)$. Hence $\beta=\theta/k$ is the contract that maximizes the principal’s utility. ∎
3.2 Discussion of the Result
The uniformity result is particularly attractive from a practical perspective. It implies that the principal can achieve her maximum utility and ensure fairness simultaneously. In particular, there is no need to discriminate against agents whose productivity may differ (as long as their return to scale parameter $k$ remains the same). This standardization is particularly relevant in industries like franchising or E-commerce, where the principal (the franchiser or the E-commerce platform) contracts with multiple agents (franchisees, E-commerce workers / content creators), and profit must be shared between the two parties. From a practical perspective, using uniform contracts streamlines the creation of contracts and also promotes trust between principals and agents.
As a special case, the results explain the widespread 50-50 split rule seen in many scenarios (e.g., [22]). Specifically, when the agents have a quadratic cost function ($k=2$), the results indicate that the 50-50 split is precisely the optimal linear contract which ensures maximum utility for the principal.
4 Estimating the Optimal Contract: an Instrumental Regression Approach
Figure 1: Causal relationship between variables. The agent’s response is shaded indicating it is unobserved by the principal. The dashed item is a second set of observed signals, corresponding the the scenario studied in Section 5 when repeated observations are available.
Algorithm 1 Use Contract as IV: Explore then Commit Algorithm
[…]
for $t=1,\dots,\tau$ do
     // If using basis vector for ‘exploration’:
     Set $\beta_{t}=\mathbf{e}_{i}$, here $i=\lceil t/\sqrt{T}\rceil$
     // If using small random perturbation for ‘exploration’:
     Set $\beta_{t}=(\theta_{0}+\rho_{t})/k$, here $\rho_{t}$ is a suitable random vector with variance $\sigma_{0}^{2}$ in each coordinate
     Observe the signal $x_{t}\in\mathbb{R}^{d}$ and the total revenue $y_{t}\in\mathbb{R}$
end for
Record the data so far into $B_{\tau},X_{\tau},Y_{\tau}$
Use the GMM estimator and compute the estimate :
$\hat{\theta}=(B_{\tau}^{\top}X_{\tau})^{-1}B_{\tau}Y_{\tau}.$
for $t=\tau+1,\dots,T$ do
     Set $\beta_{t}=\hat{\theta}/k$ as the contract
end for
I now study a setting where the parameter $\theta$ is unknown, i.e., the marginal utility of each task is unknown. Stated a bit informally, the principal is unsure about the exact “importance” of each task. For example, when the manager (the agent) manages a firm and participates in multiple tasks, the firm’s owner (the principal) may be unsure of the exact importance of each task and how much each task contributes to the firm’s total revenue. When the parameter $\theta$ is unknown, the principal must estimate or learn this parameter through observed data. By the previous result (Theorem 1), which states the optimal contract is $\theta/k$, finding an approximately optimal contract is essentially equivalent to obtaining an accurate estimate for $\theta$. To make matters a bit simpler, assume the return to scale parameter $k$ is known and treated as a constant.
The data consists of tuples of the form $(\beta_{t},w_{t},x_{t},y_{t})$, representing the contract, agent’s true effort, observed signal, and total production outcome respectively. In the following, the total production outcome $y_{t}$ will just be referred to as the total revenue for short. Note that the observed data by the principal only include $(\beta_{t},x_{t},y_{t})$, and the agent’s true effort $w_{t}$ is hidden.
Similar to the previous section, it is assumed the observed signals for each task is unbiased:
$\mathbb{E}[x_{t}|w_{t}]=w_{t}$
and that the total revenue satisfies a linear relation:
$\mathbb{E}[y_{t}|w_{t}]=\langle\theta,w_{t}\rangle.$
Fix any agent’s type (i.e., a cost function that satisfies 1). Define $u(\beta)$ as the expected utility when posting contract $\beta$. The principal is interested in finding a contract $\beta$ that minimizes the utility loss, which is the expected utility difference between posting the optimal contract $\theta/k$ and the contract $\beta$:
$u(\theta/k)-u(\beta).$
The following two assumptions are imposed, which allows us to relate the utility loss to the squared estimation error.
Assumption 2.
$c_{0}\leq\theta_{i}\leq C_{0}$ for $i\in[d]$.
The above assumption is natural and relatively mild, it states that there is some lower and upper-bound on the marginal utility of each task.
Assumption 3.
There exists some constant $c$ so that for any $\hat{\theta}\in[c_{0},C_{0}]^{n}$, the following holds
|$\absolutevalue{u(\theta/k)-u(\hat{\theta}/k)}\leq c\cdot\norm{\theta/k-\hat{\theta}/k}_{2}^{2}.$
The above assumption can essentially be seen as a smoothness assumption on the function $u$. Then, if the principal has an estimate $\hat{\theta}$ and the principal posts $\beta=\hat{\theta}/k$ as the contract, the utility loss can be bounded by the square estimation error (up to constant factors):
|$\absolutevalue{u(\theta/k)-u(\hat{\theta}/k)}\leq O\left(\norm{\theta-\hat{\theta}}_{2}^{2}\right).$
The above will hold for any agent type as long as the assumptions are met. Therefore, the following will mainly focus on obtaining upper bounds on the quantity $\norm{\theta-\hat{\theta}}_{2}^{2}$.
The problem of finding the optimal contract can either take place in an offline or online setting. In the offline setting, a sequence of data consisting of observations $(\beta_{t},x_{t},y_{t})$ is given to the principal, and the principal must obtain an accurate estimate of the true parameter $\theta$.
In the online setting, the principal interacts with agents in a sequential fashion consisting of $T$ rounds. In this scenario, for each round $t$, the principal must choose the contract $\beta_{t}$ based on historical data, then observe the signals $x_{t}$ and the total revenue $y_{t}$. In this online setting, the principal may be interested in minimizing the cumulative utility loss. A difference from the offline setting is that the principal needs to engage in some ‘experimentation’. Specifically, the principal may want to strategically choose some contracts $\beta$ in order to get a better estimate of the true parameter $\theta$; however, this may come at the cost of losing immediate revenue.
4.1 Offline Setting: Estimation with Instrumental Regression
The causal relationship between $(\beta_{t},w_{t},x_{t},y_{t})$ is summarized in Figure 1. The problem of estimating $\theta_{t}$ is essentially a (linear) regression problem. It can be seen that the dependence between the variables form a measurement error model, and it is well-known that directly regressing on $x_{t}$ as an attempt to learn $\theta$ will lead to a biased estimate. However, it can be observed that the contract $\beta_{t}$ acts as a valid instrumental variable. Specifically, it can be observed that
$\mathbb{E}[\beta_{t}(y_{t}-\langle\theta,x_{t}\rangle)]=\mathbf{0}\ (\in\mathbb{R}^{d}).$
Therefore an instrumental regression approach and the generalized method of moments estimator can be applied. When given a sequence of offline data, the generalized method of moments estimator takes the following form:
$\hat{\theta}_{T}=(B^{\top}_{T}X_{T})^{-1}B^{\top}_{T}Y_{T}.$
Here $B_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the posted contracts each round; $X_{T}\in\mathbb{R}^{T\times d}$ is a matrix with rows representing the signals each round; $Y_{T}\in\mathbb{R}^{T}$ is a column vector representing the outcomes each round. Denote $\gamma_{t}=y_{t}-\langle\theta,x_{t}\rangle$, and collect values for $\gamma_{1:t-1}$ into the column vector $\Gamma_{t}$.
Proposition 1.
The estimation error
$\norm{\theta-\hat{\theta}_{T}}_{2}\leq\frac{\norm{B_{T}^{\top}\Gamma_{T}}}{\sigma_{\min}(B_{T}^{\top}X_{T})}.$
Proof.
$\hat{\theta}_{T}$ $=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}Y_{T}$
$=(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}(X_{T}\theta+\Gamma_{T})$
$=\theta+(B_{T}^{\top}X_{T})^{-1}B_{T}^{\top}\Gamma_{T}.$
The proposition then follows. ∎
Remark 1.
By a standard concentration inequality, the numerator can be upper bounded as $O(\sqrt{d{T}})$ with high probability. The estimation error then depends crucially on the minimum singular value of $B_{T}^{\top}X_{T}$.
4.2 Online Setting: Learning the Optimal Contract
The estimator above gives a estimation error bound when offline observations are available. However, it does not directly give rise to a learning algorithm in the online setting, where the principal needs to choose the contract each round. To give some intuition, in this setting, the principal essentially faces an optimal design problem with an exploration-exploitation tradeoff. Specifically, the matrix $B_{t}$, consisting of contracts $\beta_{t}$, is the design, and the principal must choose this matrix such that the estimate $\hat{\theta}$ becomes accurate. At the same time, the principal must ensure that the chosen contracts $\beta_{t}$ are close to the true optimal contract $\theta/k$ in order to guarantee a small cumulative utility loss.
Lower bound on regret: $\Omega(d\sqrt{T})$
Consider a setting where each agent has cost function $c(w)=\frac{1}{2}\norm{w}_{2}^{2}$, and that this cost function is known to the principal. Then, the agent’s best response under the contract $\beta$ is exactly $w=\beta$. The principal’s expected utility when a contract $\beta$ is posted is then
$u(\beta)=\langle\theta,\beta\rangle-\norm{\beta}_{2}^{2}.$
The optimal contract is then $\beta=\theta/2$ (this is also implied by the previous uniformity result). The principal now essentially faces a zero-order optimization problem with a quadratic objective function. The contract $\beta$ is the decision variable, and the realized total revenue is the observed variable. Further, by a result in [23], the tight regret bound for this problem is $\Omega(d\sqrt{T})$.
Simple algorithms achieving $O(d\sqrt{T})$ cumulative utility loss
Perhaps a bit surprisingly, a very simple algorithm can achieve $O(d\sqrt{T})$ regret. In the first $d\sqrt{T}$ rounds, for each $i\in[d]$, the principal should set the contract as $\beta=\mathbf{e}_{i}$ for $\sqrt{T}$ rounds. The agents best response $w\in\mathbb{R}^{d}$ should only have a non-zero value in the $i$-th component $w_{i}$. Hence, the data in these rounds can be used to estimate $\theta_{i}$, and the estimation error after the initial $d\sqrt{T}$ rounds is:
|$\absolutevalue{\hat{\theta}_{i}-\theta_{i}}=O(1/{T^{1/4}}).$
In the remaining rounds, the principal can simply use the estimate after the initial $d\sqrt{T}$ rounds, and post the contract as $\beta=\hat{\theta}/k$. The regret in these rounds can be bounded as:
$T\cdot\norm{\hat{\theta}-\theta}_{2}^{2}=d\sqrt{T}.$
[…]
5 Improving the Estimation with Repeated Observations and Diversity
Algorithm 2 Use Repeated Observation as IV: Pure Exploitation with Diversity
Parameters include time horizon $T$, number of tasks $d$, minimum eigenvalue $\lambda_{0}$, failure probability $\delta$
Epoch $e$ contains |$\absolutevalue{\tau_{e}}=\max\{d\cdot 2^{e},8dC_{0}K_{0}\ln(d/\delta)/(c_{0}^{2}\lambda_{0}),4\sigma_{0}^{2}d^{2}\ln(d^{2}/\delta)/(c_{0}^{2}\lambda_{0})\}$ rounds
In the 0-th epoch with $d$ rounds, post $\beta=\mathbf{e}_{i}$ for each $i\in[d]$ once and record $\widetilde{X}_{1},X_{1},Y_{1}$
for epoch $e=1,2,\cdots,\lceil\log(T/d)\rceil$ do
     Compute $\hat{\theta}_{e}$ using the GMM estimator from data in previous epoch:
$\hat{\theta_{e}}=(\widetilde{X}_{e}^{\top}X_{e})^{-1}X_{e}Y_{e}$
     Post $\beta_{t}=\hat{\theta}_{t}/2$ for each round in this epoch
     Record the data from this epoch into $\widetilde{X}_{e+1},X_{e+1},Y_{e+1}$
end for
This section continues the discussion where the principal must learn $\theta$ using sequential observations. The previous section showed that using an instrumental regression approach where the contract is the instrumental variable, the principal must balance exploration (ensuring that $\sigma_{\min}(B_{t}^{\top}X_{t})$ grows as $t$ increases) and exploitation (ensuring $\beta_{t}$ is close to the true optimal contract). In this section, I show that, under some additional assumptions, the learning rate of the principal can be greatly improved. Informally, I show that when the agents are sufficiently ‘diverse’, and that when repeated observations are available, the principal can achieve a logarithmic regret, even if the principal is using a ‘pure exploitation algorithm’.
To state the results in a more concise manner, this section restricts attention to the case where the cost functions of the agents have diagonal quadratic forms. Specifically, the cost function of the agent is parameterized by a vector $\kappa\in(\mathbb{R}^{d})^{+}$, so that the cost function of the agent is $c(w)=\frac{1}{2}w^{\top}\mathtt{diag}(\kappa)^{-1}w$. Then, the agent’s best response to a contract $\beta$ is the maximizer of $\langle\beta,w\rangle-c(w)$, which is equal to
$w=\kappa\odot\beta.$
I assume that the agent cost function in each round, parameterized by $\kappa_{t}$, is drawn from an unknown distribution. The distribution satisfies the following.
Assumption 4.
The minimum eigenvalue of $\mathbb{E}_{\kappa\sim\mathcal{D}}[\kappa\kappa^{\top}]$ is lower bounded by $\lambda_{0}$. Further, for any $\kappa$ in the support of $\mathcal{D}$, $\kappa\leq K_{0}$.
The key part in the above assumption is the minimum eigenvalue $\lambda_{0}$, it suggests that the agents’ are sufficiently “diverse" in their types. The second assumption is that repeated observations are available. I assume that for each agent, the principal can observe two sets of signals, denoted by $x_{t}$ and $\widetilde{x}_{t}$. The signals are $\sigma_{0}^{2}$-subgaussian vectors and satisfy the following:
$\mathbb{E}[x_{t}|w_{t}]=\mathbb{E}[\widetilde{x}_{t}|w_{t}]=w_{t}.$
Assumption 5.
The two sets of signals are assumed to be (conditionally) independent. Specifically:
$x_{t}\perp\widetilde{x}_{t}|w_{t}.$
The causal relationship is summarized in Figure 1 with the dashed items included. It can be observed that the second set of observations can serve as an instrumental variable for the other, specifically:
$\mathbb{E}[\widetilde{x}_{t}(y_{t}-\langle x_{t},\theta\rangle)]=\mathbf{0}\ (\in\mathbb{R}^{d}).$
Therefore, we can use the following GMM estimator to obtain an estimate $\hat{\theta}$:
$\hat{\theta}_{t}=(\widetilde{X}_{t}^{\top}X_{t})^{-1}\widetilde{X}_{t}^{\top}Y_{t}.$
The algorithm will simply use the estimator above, without regard to ‘exploration’. In the implementation, the algorithm proceeds in epochs with epochs doubling in length to save computation. In each epoch $e$, the principal sets the contract using the estimation $\hat{\theta}_{e}$ according to the above formula using data from the previous epochs. See Algorithm 2 for details.
Different from the algorithms in the previous section, the algorithm does not face the exploration-exploitation tradeoff, as the ‘diversity’ of the agents already acts as a form of exploration.
Theorem 2.
With probability $1-1/T$, the cumulative regret of the algorithm is $\widetilde{O}(d/\lambda_{0}^{2})$.
Proof Sketch. .
Consider an epoch $e$ with length |$\absolutevalue{\tau_{e}}$ |. The minimum singular value of $(\widetilde{X}_{e}^{\top}X_{e})$ is on the order of $\Omega(\absolutevalue{\tau_{e}}\lambda_{0})$. At the same time, the term $\norm{\widetilde{X}_{e}^{\top}Y_{e}}$ can be upper bounded by $\widetilde{O}(\sqrt{d\absolutevalue{\tau_{e}}})$. Hence, for each round in the epoch, the squared estimation error can be upper bounded by:
$\norm{\hat{\theta}_{e}-\theta}_{2}^{2}\leq\widetilde{O}(d/(\absolutevalue{\tau_{e}}\lambda_{0}^{2})).$
Therefore, the regret in each epoch is on the order of $\widetilde{O}(d/\lambda_{0}^{2})$. There are a total of $O(\log T)$ epochs, therefore the total regret is $\widetilde{O}(d/\lambda_{0}^{2})$. ∎
Remark 2.
At a technical level, the analysis is a combination of the convergence analysis of GMM estimators (e.g. Chapter 18 in [5]) and smoothed analysis in bandit problems ([19, 20, 21]).
Remark 3.
Interestingly, the “fast" convergence also holds under a different notion of ‘diversity’. Instead of assuming agents’ cost functions are diverse, consider a setting where agents do not necessarily best respond, and their response $w_{t}$ may be a random vector with some variance lower bounded from below. In this setting, the minimum singular value $\widetilde{X}_{t}^{\top}X_{t}$ still grows linearly with $t$, and the “fast" convergence rate is achieved.
6 Numerical Experiments
Figure 2: (a): Total loss using the Contract as IV, and exploration is achieved through small random perturbations; (b): Total loss using the repeated observations as IV and when the “diversity" condition is satisfied.
This section describes numerical simulations that test the performance of the learning algorithms based on the GMM estimators. In the experiments, the number of tasks is chosen as $d=5$, the total number of rounds is chosen as $T=10^{6}$, and the true parameter $\theta$ is chosen so that each entry is between $[1,3]$. Experiments are ran on a personal laptop.
The experiments have two settings. In the first setting, the contract is used as the instrumental variable, and small random perturbations are added as a means to achieve exploration in the exploration phase (Algorithm 1). The small perturbations are chosen as truncated Gaussian vectors so that the posted contract has components within $[1,3]$. The agents have cost function $c(w)=\norm{w}_{2}^{2}$. The cumulative utility loss of a single run is plotted in Figure 2(a). It can be seen the loss grows quickly in the warm start phase (which consists of $d\sqrt{T}\approx 5\times 10^{3}$ rounds), and then the regret grows slowly in the remaining rounds. The experiments are run for 20 trials, and the average cumulative loss in the 20 trials is 4134.7, roughly on the order of $O(d\sqrt{T})$.
In the second setting, repeated observations are used as the instrumental variable (Algorithm 2). The agent’s cost function is chosen as a diagonal quadratic function with form
$c(w)=\sum_{i=1}^{d}\frac{w_{i}^{2}}{2k_{i}}.$
The parameter $k_{i}$ is sampled from the uniform distribution on $[1,2]$ so that the ‘diversity’ condition is satisfied. The cumulative loss of a single run is plotted in Figure 2(b). The loss grows quickly in the 1st epoch (consisting of roughly $\widetilde{O}(d)$ rounds) and then grows very slowly in the remaining rounds. The experiment is repeated for 20 trials, and the average cumulative loss is 89.6 (recall the theoretical limit is $\widetilde{O}(d)$).
7 Conclusion and Limitation
This work studied the multitask principal-agent problem. I first present a result justifying the use of uniform contracts when the principal contracts with heterogeneous agents. Then, I study the problem of learning the optimal contract and identify the problem as a measurement error problem, and propose using instrumental regression with the generalized method of moments estimator to estimate and learn the optimal contract. A limitation of the current work is that in the “diversity” condition in Section 5, the agents are limited to have quadratic cost functions. I believe this can be extended to more general cost functions (beyond quadratic functions and beyond $k=2$). Specifically, the “diversity” condition will hold whenever there are some variations in the agent’s “abilities" in each task. I will leave a more formal treatment to future work.
Appendix A Proof of Theorem 2
Recall the estimator in each epoch
$\hat{\theta}_{e}=\left(\widetilde{X}_{e}^{\top}X_{e}\right)^{-1}\widetilde{X}_{e}Y_{e}.$
Denote $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and collect $\gamma_{t}$ into the column vector $\Gamma_{e}$.
Proposition 2.
$\norm{\hat{\theta}_{e}-\theta}_{2}\leq\frac{\norm{\widetilde{X}_{e}\Gamma_{e}}}{\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})}$.
To bound the estimation error, the following will lower bound $\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})$ and upper bound $\norm{\widetilde{X}_{e}\Gamma_{e}}$ separately.
[…]
Lemma 1 (Matrix Chernoff, Tropp).
[…]
$\lambda_{\min}X_{k}\geq 0,\lambda_{\max}X_{k}\leq R.$
Define
$\mu_{\min}:=\lambda_{\min}(\sum_{k}\mathbb{E}X_{k}),$
$\mu_{\max}:=\lambda_{\max}(\sum_{k}\mathbb{E}X_{k}).$
Then
$\Pr[\lambda_{\min}(\sum_{k}X_{k})\leq(1-\delta)\mu_{\min}]\leq d\cdot\left[\frac{e^{-\delta}}{(1-\delta)^{1-\delta}}\right]^{\mu_{\min}/R}.$
The following Corollary may be easier to use than the above Lemma.
Corollary 1.
Using the same setup as above, then
$\Pr[\lambda_{\min}(\sum_{k}X_{k})\leq(1-\delta)\mu_{\min}]\leq d\exp(-\delta^{2}\mu_{\min}/(2R)).$
Lemma 2.
With high probability $\delta$, $\lambda_{\min}(\sum_{t\in\tau_{e}}w_{t}w_{t}^{\top})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2$.
Proof.
[…]
$\mathbb{E}[W_{t}]$ $=\mathbb{E}[w_{t}w_{t}^{\top}]$
$=\mathbb{E}[\mathtt{diag}(\kappa_{t})\theta_{t}\theta_{t}^{\top}\mathtt{diag}(\kappa_{t})]$
$=\mathbb{E}[\Theta_{t}\kappa_{t}\kappa_{t}^{\top}\Theta_{t}]$
$=\Theta_{t}\mathbb{E}[\kappa_{t}\kappa_{t}^{\top}]\Theta_{t}$
Recall $\lambda_{\min}(\mathbb{E}[\kappa_{t}\kappa_{t}^{\top}])\geq\lambda_{0}$. The following holds
$\lambda_{\min}(\mathbb{E}W_{t})\geq c_{0}^{2}\lambda_{0}.$
By Tropp’s lemma,
$\Pr[\lambda_{\min}(\sum_{k}W_{k})\leq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2]$ $\leq d\cdot\exp(-\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/(8dC_{0}K_{0}))$
$\leq\delta$
when
|$\absolutevalue{\tau_{e}}\geq\frac{8dC_{0}K_{0}}{c_{0}^{2}\lambda_{0}}\cdot\ln(d/\delta)$
Lemma 3.
With probability $1-2\delta$, for each epoch $e$ with length as specified in Algorithm 2, $\sigma_{\min}(\widetilde{X}_{e}^{\top}X_{e})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/4$.
Proof.
In the following denote $\varepsilon_{t}=x_{t}-w_{t},\widetilde{\varepsilon}_{t}=\widetilde{x}_{t}-w_{t}$.
$\widetilde{X}_{e}^{\top}X_{e}$ $=\sum_{t\in\tau_{e}}x_{t}\widetilde{x}_{t}^{\top}$
$=\sum_{t\in\tau_{e}}(w_{t}+\varepsilon_{t})(w_{t}+\widetilde{\varepsilon}_{t})^{\top}$
$=\sum_{t\in\tau_{e}}w_{t}^{\top}w_{t}+\varepsilon_{t}^{\top}w_{t}+w_{t}\varepsilon_{t}+\varepsilon_{t}^{\top}\widetilde{\varepsilon}_{t}$
The previous lemma showed that with probability $1-\delta$,
$\lambda_{\min}(\sum_{t\in\tau_{e}}w_{t}^{\top}w_{t})\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/2.$
The remaining term
$\sum_{t\in\tau_{e}}\varepsilon_{t}w_{t}^{\top}+w_{t}\widetilde{\varepsilon}_{t}^{\top}+\varepsilon\widetilde{\varepsilon}^{\top}$
is a random matrix with mean $\mathbf{0}$. By a standard concentration inequality, the absolute value of each entry in the matrix can be upper-bounded as |$\sqrt{\absolutevalue{\tau_{e}}\sigma_{0}^{2}\ln(d^{2}/\delta)}$ with probability $\delta$. Then, since the minimum singular value must be upper bounded by the Frobenius norm:
$\sigma_{\min}\left(\sum_{t\in\tau_{e}}\varepsilon_{t}w_{t}^{\top}+w_{t}\widetilde{\varepsilon}_{t}^{\top}+\varepsilon\widetilde{\varepsilon}^{\top}\right)\leq\sqrt{\absolutevalue{\tau_{e}}\sigma_{0}^{2}d^{2}\ln(d^{2}/\delta)}.$
The conclusion is that with probability $1-2\delta$,
[…]
|$\geq\absolutevalue{\tau_{e}}c_{0}^{2}\lambda_{0}/4.$
[…]
add sometinghere, condition on the length of $\tau_{e}$ This finishes the proof. ∎
Lemma 4.
[…]
Proof.
Recall $\gamma_{t}=y_{t}-\langle x_{t},\theta\rangle$, and $\Gamma_{e}$ is the column vector collecting the variables $\gamma_{t}$. Note that $\gamma_{t}\widetilde{x}_{t}$ is a zero-mean, $\sigma_{0}C_{0}K_{0}$-subgaussian random vector and independent conditioned on all previous steps. The result then follows from a standard concentration inequality. ∎
Theorem 3.
With probability $1-1/T$, the total regret is $\widetilde{O}(d/\lambda_{0}^{2})$.
Proof.
By Lemma 4, Lemma 3, and Proposition 2, with probability $1-3\delta$, the regret in each epoch can be upper bounded as |$\absolutevalue{\tau_{e}}\cdot\widetilde{O}(d/(\absolutevalue{\tau_{e}}\lambda_{0}^{2}))=\widetilde{O}(d/\lambda_{0}^{2})$. Choose some appropriate $\delta$, e.g., $\delta<O(1/T^{2})$. Since there are at most $\log T$ epochs, the total regret will be $\widetilde{O}(d/\lambda_{0}^{2})$. ∎